MASS-ANALYZED THRESHOLD IONIZATION OF LANTHANUM OXIDE CLUSTERS: La_{2}O_{2} AND La_{3}O_{4}

\underline{LU WU}, CHANGHUA ZHANG, SERGIY KRASNOKUTSKI, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Lanthanum oxide clusters are produced by laser vaporization in a pulsed cluster source and identified by photoionization mass spectrometry. Vibrationally resolved ion spectra are obtained with mass-analyzed threshold ionization (MATI) spectroscopy. The MATI spectra of La_{2}O_{2} and La_{3}O_{4} exhibit a very strong 0-0 transition, indicating similar geometries for the neutral and ionized clusters and a very weakly bonding or non-bonding electron ejected from an outmost molecular orbital. The ionization energies of La_{2}O_{2} and La_{3}O_{4} are measured to be 36937(5) and 28028(5) cm^{−1}, respectively. In addition, the spectra of both clusters display a number of vibrational intervals that are associated with metal-metal, metal-oxygen, and oxygen-oxygen vibrations. Preliminary data analysis shows that the La_{2}O_{2} cluster has a D_{2h} planar structure and La_{3}O_{4} has a C_{3v} cage-like structure, both with alternating La-O-La bonds. The spectra may be assigned to the ^{2}Ag←^{1}Ag transition in the case of La_{2}O_{2} and ^{1}A_{1}←^{2}A_{1} in La_{3}O_{4} .